If you push a crate with 40 N of force at an angle of 30o and your friend pushes with a force of 40 N horizontally, the amount of work you are doing is _______________ the amount of work your friend is doing on the crate.
1
Expert's answer
2020-11-02T09:23:12-0500

If you push a crate with 40 N of force at an angle of 30° and your friend pushes with a force of 40 N horizontally, the amount of work you are doing is less than the amount of work your friend is doing on the crate.


Why?


The work is


"W=Fd\text{ cos}\alpha."

The angle alpha here is the angle between the force and displacement vectors. So, the greatest value is when "\text{cos}\alpha=1, \alpha=0°."